# Einstein connection of nonsymmetric Riemannian manifold

### 04 Mar 2026 v2

@@ 4 Einstein connection of an NGT space
where $\delta_{4}(X,Y,Z)$ is given in Section 5. Using (2.8) and the $f^{2}$-torsion condition, we reduce $\delta_{4}(X,Y,Z)$ to $\delta_{5}(X,Y,Z)$, given in ection 5. Using (4.1) and (4.1), we find
$T(fY,fZ,(3\,\widetilde{Q}$ $+2\,\widetilde{Q}^{2})X)=T(Y,Z,(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X)+dF(Y,Z,(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X)$
$-dF(fY,fZ,(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})X)-2\,(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,Y,fZ)$
$+2\,(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,fY,Z),$
$T(Y,fZ,\widetilde{Q}f^{3}X)$ $+T(fY,Z,\widetilde{Q}f^{3}X)=dF(Y,Z,\widetilde{Q}f^{2}({I}+f^{2})X)+2\,(\nabla^{g}F)(\widetilde{Q}f^{2}X,Y,Z)$
$+2\,(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,Y,fZ)-2\,(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,fY,Z),$
and then rewrite $\delta_{5}(X,Y,Z)$ as
$\delta_{5}(X,Y,Z)=$ $\,-T(Y,Z,(9\,\widetilde{Q}+7\,\widetilde{Q}^{2}+2\,\widetilde{Q}^{3})X)+T(Y,Z,(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X)$
$+dF(Y,Z,(3\,\widetilde{Q}{+}2\,\widetilde{Q}^{2})f^{2}X){-}dF(Y,Z,\widetilde{Q}f^{2}({I}+f^{2})X){+}dF(Y,Z,\widetilde{Q}f^{2}X)$
$-2\,dF(fY,fZ,(\widetilde{Q}+\widetilde{Q}^{2})X)-2\,(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,Y,fZ)$
$+2\,(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,fY,Z)-2\,(\nabla^{g}F)(\widetilde{Q}f^{2}X,Y,Z)$
$-2\,(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,Y,fZ)+2\,(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,fY,Z)$
$-dF(Y,fZ,\widetilde{Q}f^{3}X)-dF(fY,Z,\widetilde{Q}f^{3}X).$
Note that $\delta_{i}(X,Y,Z)\ (1\leq i\leq 5)$ vanish when $\widetilde{Q}=0$. From the equation above and (4) with $\delta_{4}(X,Y,Z)$ replaced by $\delta_{5}(X,Y,Z)$, the equation (4.3) follows. If $fX=0$, then from (4.1) we directly get (4.5). ∎
@@
Take almost Hermitian manifolds $(M_{j},J_{j},g_{j})\ (j=1,\dots,m)$, where $J_{j}^{2}=-{I}_{\,j}$. Consider the Riemannian product $(M,g)=\bigl(M_{1}\times\cdots\times M_{m},\;g_{1}\oplus\cdots\oplus g_{m}\bigr)$, and define the (1,1)(1,1)–tensor $f=\sqrt{\lambda_{1}}\,J_{1}\oplus\cdots\oplus\sqrt{\lambda_{m}}\,J_{m}$. Then $(M,f,g)$ is a weak almost Hermitian manifold with $f^{2}=-\bigoplus_{j}\lambda_{j}\,{I}_{j}$, and on $TM_{j}$ (and for $X\in TM_{j}$) we have
@@
Since the Levi-Civita connection of the Riemannian product preserves the splitting $TM=\bigoplus_{j=1}^{m}TM_{j}$, $f^{2}$ acts as a multiple of the identity on each factor. The 2-form $F$ corresponding to $f$ is given by $F=F_{1}\oplus\cdots\oplus F_{m}$, where $F_{j}(X,Y)=g_{j}(X,\sqrt{\lambda}_{j}J_{j}Y)$. As a consequence,
@@
If an Einstein connection $\nabla$ satisfies the $f^{2}$-torsion condition (3.3), then $T(X,Y)=0$ for $X\in TM_{i},\,Y\in TM_{j}\ (i\neq j)$. Hence, the torsion tensor splits into the direct sum $T=T|_{TM_{1}}\oplus\cdots\oplus T|_{TM_{m}}$. Substituting $\widetilde{Q}|_{\,TM_{j}}=(\lambda_{j}-1)\,{I}_{j}$ into (4.3), then using
@@
$\ -\tfrac{1}{4}\,\lambda_{j}^{-2}(\lambda_{j}-1)(3\lambda_{j}+1)dF_{j}(Y,Z,X)+\tfrac{1}{4}(\lambda_{j}-1)\lambda_{j}^{-1/2}dF_{j}(Y,J_{j}Z,J_{j}X)$
$\ +\tfrac{1}{4}(\lambda_{j}-1)\lambda_{j}^{-1/2}dF_{j}(J_{j}Y,Z,J_{j}X)-\tfrac{1}{2}\lambda_{j}^{-2}(\lambda_{j}-1)dF_{j}(J_{j}Y,J_{j}Z,X).$
If each $(M_{j},J_{j},g_{j})$ is a Kähler manifold, then $\nabla^{g_{j}}F_{j}=0$ and $dF_{j}=0$. In this case, all torsion components vanish, and the Einstein connection reduces to the Levi-Civita connection. Using (2.5), we can further simplify the formula for $T|_{TM_{j}}$ as
$T|_{TM_{j}}(Y,Z,X)=$ $\ \frac{(\lambda-1)(5-3\lambda)}{4\lambda^{2}}\,(\nabla^{g_{j}}_{X}F_{j})(Y,Z)+\Bigl(\frac{1}{\lambda}+\frac{\lambda-1}{2\sqrt{\lambda}}\Bigr)\,(\nabla^{g_{j}}_{J_{j}X}F_{j})(Y,J_{j}Z)$
$-\Bigl(\frac{3\lambda^{2}-2\lambda+1}{4\lambda^{2}}+\frac{\lambda-1}{4\sqrt{\lambda}}\Bigr)\,\big\{(\nabla^{g_{j}}_{Z}F_{j})(X,Y)+(\nabla^{g_{j}}_{Y}F_{j})(Z,X)\big\}$
$+\Bigl(\frac{\lambda-1}{4\sqrt{\lambda}}{-}\frac{1}{2\lambda}{-}\frac{\lambda-1}{2\lambda^{2}}\Bigr)\,\big\{(\nabla^{g_{j}}_{J_{j}Y}F_{j})(J_{j}Z,X)+(\nabla^{g_{j}}_{J_{j}Z}F_{j})(X,J_{j}Y)\big\}.$
The formula above for $\lambda=1$ gives the solution (2.8).
@@
Let $\nabla$ be an Einstein connection of a weak almost Hermitian manifold $(M^{2n},f,g>0)$, considered as a nonsymmetric Riemannian space $(M,G=g+F)$ with $F(X,Y)=g(X,fY)$. If the condition (3.3) is valid, then the torsion $T$ of $\nabla$ is given by (4.3).
@@
For a nonsymmetric Riemannian space $(M,G=g+F)$, the tensor $P={I}-f^{2}$ is positive definite, and therefore non-degenerate. Using (2.5), we can further simplify (4.3) to the form with $\nabla^{g}\,F$-terms.
Corollary 4.7 (see [12]).
Let $(M^{2n},f,g)$ be a weak almost Hermitian manifold with fundamental $2$-form $F$, considered as an NGT space $(M,G=g+F,\nabla)$, and $\nabla$ (an Einstein connection) has a totally skew-symmetric torsion $(0,3)$-tensor $T$. If the $f$-torsion condition (2.17) is true, then $(M^{2n},f,g)$ is a weak nearly Kähler manifold and the following is true:
$T(fX,Y,Z)=-\tfrac{1}{3}\,dF(fX,Y,Z),\quad\nabla\,f^{2}=\nabla^{g}f^{2}=0.$
@@ 5 The terms $\delta_{i}(X,Y,Z)$
Here, we present the terms $\delta_{i}(X,Y,Z)\ (1\leq i\leq 5)$, see Section 4. Their expressions were verified with the help of MAPLE program.
@@
$-T(Z,X,\widetilde{Q}Y)-T(X,Y,\widetilde{Q}Z)+T(\widetilde{Q}Y,fZ,fX)-T(fX,Y,\widetilde{Q}fZ)$
$-T(Y,\widetilde{Q}Z,X)-T(Y,\widetilde{Q}fZ,fX)-T(Y,Z,\widetilde{Q}X)+T(\widetilde{Q}fZ,X,fY)$
$-T(Z,\widetilde{Q}X,Y)+T(\widetilde{Q}fX,Y,fZ)-T(fY,Z,\widetilde{Q}fX)-T(Z,\widetilde{Q}fX,fY)$
$+T(\widetilde{Q}Z,fX,fY)+T(\widetilde{Q}X,fY,fZ).$
@@
$-T(Y,fZ,\widetilde{Q}f^{3}X)-T(fY,Z,\widetilde{Q}f^{3}X)-dF(Y,fZ,\widetilde{Q}f^{3}X)$
$-dF(fY,Z,\widetilde{Q}f^{3}X)+dF(fY,fZ,\widetilde{Q}X)+dF(Y,Z,\widetilde{Q}f^{2}X).$
@@ Conclusion
We presented explicitly the Einstein connection for nonsymmetric pseudo-Riemannian, in particular, for weak almost Hermitian manifolds, satisfying the $f^{2}$-torsion condition. We also obtained the corresponding formulas for almost contact metric manifolds, assuming the $f^{2}$-torsion condition. We expressed the torsion in terms of $\nabla^{g}F$ and $dF$, and the additional $\widetilde{Q}-$terms, explained precisely how the weak almost Hermitian case differs from the almost Hermitian case. The presented identities provide a new tool for constructing examples and for studying classes of nonsymmetric pseudo-Riemannian manifolds, in particular, weak almost Hermitian manifolds with the usual Gray-Hervella classification, in Einstein’s nonsymmetric gravitational theory. In further research we will study weak (para-) metric $f$-manifolds, in particular, weak (para-) contact metric manifolds, with an Einstein connections or connections with totally skew-symmetric torsion, satisfying the $f^{2}$-torsion condition for vectors orthogonal to the distribution $\ker f$.
Acknowledgments. This work was partially supported by the Ministry of Education, Science and Technological Development of the Republic of Serbia (contract reg. no. 451-03-34/2026-03/200124 for Milan Zlatanović). The authors would also like to thank Miroslav Maksimović for carefully reading the paper and for his helpful and valuable comments.

### 24 Mar 2026 v3

@@ 4 Einstein connection of an NGT space
where $\delta_{4}(X,Y,Z)$ is given in Section 5. Using (2.8), $f^{2}{=}I+\widetilde{Q}$, and the $f^{2}$-torsion condition (4.1), we reduce $\delta_{4}(X,Y,Z)$ to $\delta_{5}(X,Y,Z)$, given in Section 5. Using (4.1) and (4.1), we get
$T(fY,fZ,(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})X)=T(Y,Z,(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X)+dF((3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,Y,Z)$
$\qquad-dF((3\,\widetilde{Q}+2\,\widetilde{Q}^{2})X,fY,fZ)-2\,(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,Y,fZ)$
$\qquad+2\,(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,fY,Z),$
$T(Y,fZ,\widetilde{Q}f^{3}X)+T(fY,Z,\widetilde{Q}f^{3}X)=T(Y,Z,\widetilde{Q}^{2}f^{2}X)-dF(\widetilde{Q}^{2}f^{2}X,Y,Z)$
$\qquad+2\,(\nabla^{g}F)(\widetilde{Q}f^{2}X,Y,Z)+2\,(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,Y,fZ)-2\,(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,fY,Z),$
which allow us to transform some terms in $\delta_{5}(X,Y,Z)$ and then rewrite it as
$\delta_{6}(X,Y,Z)=$ $\,-T(Y,Z,\ (12\,\widetilde{Q}+13\,\widetilde{Q}^{2}+6\,\widetilde{Q}^{3}+\widetilde{Q}^{4})X)-3\,dF(\widetilde{Q}f^{4}X,Y,Z)$
$-dF((3\,\widetilde{Q}+2\,\widetilde{Q}^{2})X,fY,fZ)-2\,(\nabla^{g}F)(P^{-1}(2\,\widetilde{Q}+\widetilde{Q}^{2})f^{2}X,Y,fZ)$
$+2\,(\nabla^{g}F)(P^{-1}(2\,\widetilde{Q}+\widetilde{Q}^{2})f^{2}X,fY,Z)-2\,(\nabla^{g}F)(\widetilde{Q}f^{2}X,Y,Z).$
Note that $\delta_{i}(X,Y,Z)\ (1\leq i\leq 6)$ vanish when $\widetilde{Q}=0$. From the equation (4) with $\delta_{4}(X,Y,Z)$ replaced by $\delta_{6}(X,Y,Z)$, the equation (4.3) follows. Using (2.5), we can further simplify (4.3) to the form with $\nabla^{g}\,F$-terms and without $dF$-terms. ∎
@@
Take almost Hermitian manifolds $(M_{j},J_{j},g_{j})\ (j=1,\dots,m)$, where $J_{j}^{2}=-{I}_{\,j}$. Consider the Riemannian product $(M,g)=\bigl(M_{1}\times\cdots\times M_{m},\;g_{1}\oplus\cdots\oplus g_{m}\bigr)$, and define the (1,1)(1,1)–tensor $f=\sqrt{\lambda_{1}}\,J_{1}\oplus\ldots\oplus\sqrt{\lambda_{m}}\,J_{m}$, where $\lambda_{j}\in\mathbb{R}_{+}$. Then $(M,f,g)$ is a weak almost Hermitian manifold with $f^{2}=-\bigoplus_{j}\lambda_{j}\,{I}_{j}$, and on $TM_{j}$ and for $X\in TM_{j}$ we get
@@
Since the Levi-Civita connection of the Riemannian product preserves the splitting $TM=\bigoplus_{j=1}^{m}TM_{j}$, $f^{2}$ acts as a multiple of the identity on each factor. The 2-form $F$ corresponding to $f$ is given by $F=F_{1}\oplus\ldots\oplus F_{m}$, where $F_{j}(X,Y)=g_{j}(X,\sqrt{\lambda}_{j}J_{j}Y)$. As a consequence,
@@
If an Einstein connection $\nabla$ satisfies the $f^{2}$-torsion condition (3.3), then $T(X,Y)=0$ for $X\in TM_{i},\,Y\in TM_{j}\ (i\neq j)$. Hence, the torsion tensor splits into the direct sum $T=T|_{TM_{1}}\oplus\ldots\oplus T|_{TM_{m}}$. Substituting $\widetilde{Q}|_{\,TM_{j}}=(\lambda_{j}-1)\,{I}_{j}$ into (4.3), then using
@@
$\ -\tfrac{1}{4}\,\lambda_{j}^{-2}(\lambda_{j}-1)(3\lambda_{j}+1)dF_{j}(X,Y,Z)+\tfrac{1}{4}(\lambda_{j}-1)\lambda_{j}^{-1/2}dF_{j}(J_{j}X,Y,J_{j}Z)$
$\ +\tfrac{1}{4}(\lambda_{j}-1)\lambda_{j}^{-1/2}dF_{j}(J_{j}X,J_{j}Y,Z)-\tfrac{1}{2}\lambda_{j}^{-2}(\lambda_{j}-1)dF_{j}(X,J_{j}Y,J_{j}Z).$
If each $(M_{j},J_{j},g_{j})$ is a Kähler manifold, then $\nabla^{g_{j}}F_{j}=0$ and $dF_{j}=0$. In this case, all torsion components vanish, and the Einstein connection reduces to the Levi-Civita connection. Using (2.5), we can further simplify the formula for $T|_{TM_{j}}\ (1\leq j\leq m)$ as
$T|_{TM_{j}}(Y,Z,X)=$ $\ \frac{(\lambda_{j}-1)(5-3\lambda_{j})}{4\lambda_{j}^{2}}\,(\nabla^{g_{j}}_{X}F_{j})(Y,Z)+\Bigl(\frac{1}{\lambda_{j}}+\frac{\lambda_{j}-1}{2\sqrt{\lambda_{j}}}\Bigr)\,(\nabla^{g_{j}}_{J_{j}X}F_{j})(Y,J_{j}Z)$
$-\Bigl(\frac{3\lambda_{j}^{2}-2\lambda_{j}+1}{4\lambda_{j}^{2}}+\frac{\lambda_{j}-1}{4\sqrt{\lambda_{j}}}\Bigr)\,\big\{(\nabla^{g_{j}}_{Z}F_{j})(X,Y)+(\nabla^{g_{j}}_{Y}F_{j})(Z,X)\big\}$
$+\Bigl(\frac{\lambda_{j}-1}{4\sqrt{\lambda_{j}}}-\frac{2\,\lambda_{j}-1}{2\lambda_{j}^{2}}\Bigr)\,\big\{(\nabla^{g_{j}}_{J_{j}Y}F_{j})(J_{j}Z,X)+(\nabla^{g_{j}}_{J_{j}Z}F_{j})(X,J_{j}Y)\big\}.$
The formula above for $\lambda_{j}=1$ gives the solution (4.6).
For a nonsymmetric Riemannian space $(M,G=g+F)$, the tensor $P={I}-f^{2}$ is positive definite, and hence, non-degenerate. Therefore, we have the following.
@@
Let $\nabla$ be an Einstein connection of a weak almost Hermitian manifold $(M^{2n},f,g>0)$, considered as a nonsymmetric Riemannian space with $F(X,Y)=g(X,fY)$. If the condition (3.3) is valid, then the torsion $T$ of $\nabla$ is given by (4.3).
Corollary 4.6 (see Theorems 1 and 2 of [8]).
Let $(M^{2n},J,g)$ be an almost Hermitian manifold considered as a nonsymmetric Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,JY)$. Then the torsion of an Einstein connection $\nabla$ on $M$ is given by
$-(\nabla^{g}_{Y}F)(Z,X)-(\nabla^{g}_{Z}F)(X,Y);$ (4.14)
in particular, the torsion of a special Einstein connection $\nabla$ on $M$, see (2.19), is given by
@@
In view of (4.6), the Einstein connection $\nabla$ on an almost Hermitian manifold is special, see definition (2.19), if and only if the following identity holds:
Example 4.8.
If $(\nabla^{g}_{X}J)X=0\ (X\in\mathfrak{X}_{M})$, or $\nabla^{g}J=0$, then a weak almost Hermitian manifold $(M^{2n},J,g)$ is a weak nearly Kähler manifold, or a weak Kähler manifold, respectively, [9]. Let $(M,f,g)$ be a weak nearly Kähler manifold, considered as a nonsymmetric Riemannian manifold $(M,G=g+F)$ with $F(X,Y)=g(X,fY)$. If $\nabla$ is an Einstein connection on $M$ with totally skew-symmetric torsion $T$, then the $f$-torsion condition holds:
$T(fX,Y)=T(X,fY)=-f\,T(X,Y),$ (4.16)
@@ 5 The terms $\delta_{i}(X,Y,Z)$
Here, we present $\delta_{i}(X,Y,Z)\ (1\leq i\leq 5)$, see Section 4, verified with the help of MAPLE.
@@
$-T(Z,X,\widetilde{Q}Y)-T(X,Y,\widetilde{Q}Z)+T(\widetilde{Q}Y,fZ,fX)-T(fX,Y,\widetilde{Q}fZ)-T(Y,\widetilde{Q}Z,X)$
$-T(Y,\widetilde{Q}fZ,fX)-T(Y,Z,\widetilde{Q}X)+T(\widetilde{Q}fZ,X,fY)-T(Z,\widetilde{Q}X,Y)+T(\widetilde{Q}fX,Y,fZ)$
$-T(fY,Z,\widetilde{Q}fX)-T(Z,\widetilde{Q}fX,fY)+T(\widetilde{Q}Z,fX,fY)+T(\widetilde{Q}X,fY,fZ).$
@@
$-T(Y,fZ,\widetilde{Q}f^{3}X)-T(fY,Z,\widetilde{Q}f^{3}X)-dF(\widetilde{Q}f^{3}X,Y,fZ)$
$-dF(\widetilde{Q}f^{3}X,fY,Z)+dF(\widetilde{Q}X,fY,fZ)+dF(\widetilde{Q}f^{2}X,Y,Z).$
@@ Conclusion
We presented explicitly the Einstein connection for nonsymmetric pseudo-Riemannian, in particular, for weak almost Hermitian manifolds, satisfying the $f^{2}$-torsion condition. We also obtained the corresponding formulas for almost contact metric manifolds, assuming the $f^{2}$-torsion condition. We expressed the torsion in terms of $\nabla^{g}F$ and $dF$, and the additional $\widetilde{Q}-$terms, explained precisely how the weak almost Hermitian case differs from the almost Hermitian case. The presented identities provide a new tool for constructing examples and for studying classes of nonsymmetric pseudo-Riemannian manifolds, in particular, weak almost Hermitian manifolds with the usual Gray-Hervella classification, in Einstein’s nonsymmetric gravitational theory. In further research we will study weak (para-) metric $f$-manifolds, in particular, weak (para-) contact metric manifolds, with an Einstein connections or connections with totally skew-symmetric torsion, satisfying the $f^{2}$-torsion condition.
Acknowledgments. This work was supported for Prof. Milan Zlatanović by the Ministry of Education, Science and Technological Development of the Republic of Serbia (contract reg. no. 451-03-34/2026-03/200124). The authors would also like to thank Dr. Miroslav Maksimović for carefully reading the paper and for his helpful and valuable comments.